OpenStudy (anonymous):

log2(8-6x) = 5 and please show work thx

OpenStudy (anonymous):

8-6x=2^5

OpenStudy (anonymous):

8 - 6x = 32 x = -4

OpenStudy (anonymous):

log2(8-6x) = 5 8+6x=2^5 6x=32-8 x=24/6=4

OpenStudy (anonymous):

thanks Nick and everyone !

OpenStudy (anonymous):

-4 <---